2kbaby24:

Simplify the following expression and write the expression in standard form. 5x+9x^2−11x+23x^2

DumbFounded:

im confused

DumbFounded:

whats with both of the 2's?

AZ:

That took a little while to decode, but to confirm is this the equation \( 5x+9x^2−11x+23x^2\)

Angle:

I agree with AZ's interpretation, the question has been edited to reflect this

QueenMo12345:

First you would combine like terms: 5x+9x^2-11x+23x^2 -6x+9x^2+23x^2 Combine like terms -6x+9x^2+23x^2 -6x+32x^2 Rearrange the terms: -6x+32x^2 32x^2-6x Solution is : 32x^2-6x

TyHawkins:

Let's simplify step-by-step. 5x+9x2−11x+23x2 =5x+9x2+−11x+23x2 Combine Like Terms: =5x+9x2+−11x+23x2 =(9x2+23x2)+(5x+−11x) =32x2+−6x Answer: =32x2−6x

TyHawkins:

we are learning this in my math class